# AlphaSyndrome: Tackling the Syndrome Measurement Circuit Scheduling Problem for QEC Codes

### v1

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 21–26, 2026; Pittsburgh, PA, USA.booktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’26), March 21–26, 2026, Pittsburgh, PA, USAisbn: 979-8-4007-2359-9/2026/03doi: 10.1145/3779212.3790123submissionid: V2fp0576ccs: Computer systems organization Quantum computingccs: Hardware Quantum error correction and fault toleranceccs: Software and its engineering Compilers
@@ 2.3. Monte Carlo Tree Search
    Selection: Starting from the initial state $S_{0}$, the algorithm traverses downward the tree by selecting child nodes that maximize the ”upper confidence bound applied to trees” (UCT) (Kocsis and Szepesvári, 2006). For the $i$th child: $UCT_{i}=\hat{E_{i}}+c\sqrt{\frac{\ln N}{n_{i}}}$ where $\hat{E_{i}}$ is the expectation score of the $i$th child, $n_{i}$ is the number of times the child has been visited, $N$ is the total number of times the parent node has been visited, and $c$ is an exploration parameter, usually $\sqrt{2}$.
@@
    Expansion: The selection is continuously applied downward to the tree until either a termination or a not fully explored (it has non-child next states) node $S_{l}$. If $S_{l}$ is a termination state, then define $S_{e}:=S_{l}$ and proceed to the next step. Otherwise, we create a new child $S_{new}$ of $S_{l}$ if $\exists m\in mov(S_{l}),S_{l}\xrightarrow{m}S_{new}$ and $S_{new}$ is not already $S_{l}$’s child. We set $S_{e}:=S_{new}$. $S_{e}$ is referred to as the ”expanded node”.
@@
    Simulation: From $S_{e}$, a ”roll-out” is performed to the termination. Specifically, we continuously select a random move starting from $S_{e}$ until we reach a termination state $S_{ter}$. An evaluation is applied to $eva(S_{ter})=E$.
@@ 5.1.1. QECCs and Decoders
We select multiple QECCs and decoders as the benchmarks, ranging across codes that encode single or multiple logical qubits. They include: rotated surface code (Bombín and Martin-Delgado, 2007; Horsman et al., 2012), color code (Bombin and Martin-Delgado, 2006; Rodriguez et al., 2024), $XZZX$ code (Bonilla Ataides et al., 2021), surface code with defects (Kitaev, 2003), bivariate bicycle (BB) code (Bravyi et al., 2024b), hyperbolic surface code (Breuckmann and Terhal, 2016; Albuquerque et al., 2009), and hyperbolic color code (Delfosse, 2013; da Silva and Jr, 2018; Vuillot and Breuckmann, 2022). For the decoder, we include the minimum-weight perfect matching (MWPM) (Higgott and Gidney, 2025), BP-OSD (Roffe et al., 2020), and hypergraph unionfind (Berlekamp et al., 1978; Lovász, 1972) decoders.
@@ 5.3.2. Space-Time Resource Estimation and Reduction
We estimate absolute execution time using the IBM Brisbane device model, where a two-qubit gate costs approximately $600$ ns and a measurement takes approximately $4000$ ns (Zheng and Kanno, 2024). Single-qubit gate latency is negligible at this scale. The execution time of one syndrome-measurement round is therefore computed as: $T_{\text{round}}=\#\text{depth}\times T_{2Q}+T_{\text{meas}}$, and the corresponding space-time volume is given by
$\text{Space-Time Volume}=T_{\text{round}}\times\#\text{physical qubits}.$
Table 3 reports both the absolute execution time and the resulting space–time volume. While AlphaSyndrome sometimes incurs a higher depth in terms of two-qubit gates, this does not significantly increase—and in some cases even reduces—the absolute execution time required to achieve a given logical error rate. This effect arises from two factors. First, to reach comparable reliability, lowest-depth baseline schedules require larger code distances, whose minimal-depth schedules themselves have larger circuit depths. Second, syndrome-measurement rounds are dominated by the ancilla measurement latency, which is substantially longer than two-qubit gate latency (this holds for most hardware platforms); as a result, moderate increases in two-qubit gate depth are largely amortized by the fixed measurement cost.
@@ 5.5. Cross-Decoder Evaluation
Table 4 shows that AlphaSyndrome tailors its schedule to the target decoder. When tested with BP‑OSD, schedules compiled using BP‑OSD reduce the overall logical error rate by 25.4% on average compared to schedules compiled with unionfind, winning in 7/8 instances and peaking at 92.4% reduction on $[[7,1,3]]$. Symmetrically, when tested with unionfind, schedules compiled using unionfind reduce errors by 34.3% on average over BP‑OSD‑compiled schedules, again in 8/8 instances, with a peak of 55.53%. Overall, cross‑testing confirms that using a schedule compiled for a different decoder generally degrades performance, indicating that AlphaSyndrome indeed adapts the measurement‑order to decoder behavior.
@@ 6. Related Work and Discussion
6.1. Syndrome Measurement Scheduling
The syndrome measurement circuit scheduling problem for QEC codes plays an essential role in realizing fault-tolerant quantum computation. Some prior works on QEC discuss circuit scheduling for different QEC codes. For example,  (Horsman et al., 2012; Tomita and Svore, 2014; Dennis et al., 2002) examined circuit scheduling for surface code. In  (Tomita and Svore, 2014), the authors propose executing the gate following an S- or Z-shape ordering to minimize the effect of hook error. However, beyond surface code, this problem has not been well optimized. For instance, (Lacroix et al., 2025) adopts a trivial ordering using the qubit index for measuring syndromes in their color code experiments. Similarly,  (Pecorari et al., 2025) also adopts the trivial ordering for their hypergraph product code experiment, though this is a result of an analysis of hook error on effective code distance on hypergraph product code (Manes and Claes, 2025). Furthermore,  (Bravyi et al., 2024a) designed the syndrome extraction circuit to have minimal circuit depth for the BB code. However, these prior works only focus on a specific family of code, and their manual design approach cannot be automatically extended to other codes. Moreover, some works  (Lacroix et al., 2025; Pecorari et al., 2025) ignore the influence of hook error and focus only on the overall depth of the circuit. Even for works that consider hook error  (Tomita and Svore, 2014; Manes and Claes, 2025), their analysis remains at the circuit level without evaluating the logical error rate. Moreover, the approximation in the decoder and the non-uniform errors are not yet considered, leading to performance degradation. In contrast, the AlphaSyndrome proposed in this paper can automatically take all these factors into consideration and yield a syndrome measurement circuit implementation with much lower logical error rates.
6.2. QEC Mapping and Gadget Synthesis
Several prior works develop synthesis techniques for other components of the QEC stack. QECC-Synth (Yin et al., 2025) employs MaxSAT to synthesize QEC layouts on sparse hardware architectures by optimizing qubit placement, routing structures, and ancilla-bridge usage, while Chiew et al. (Chiew et al., 2024) study SWAP-based embedding techniques that map fault-tolerant circuits onto constrained-connectivity hardware while preserving fault-tolerance guarantees. These approaches operate at the mapping and layout layer of the stack and do not modify the internal execution order of stabilizer measurements.
@@ 6.2. QEC Mapping and Gadget Synthesis
6.3. MCTS-Based Quantum Synthesis and Routing
MCTS has been applied in several quantum synthesis and optimization contexts, including AlphaRouter (Tang et al., 2024) and other MCTS- or reinforcement-learning–based approaches for routing and unitary synthesis (Valcarce et al., 2025; Rietsch et al., 2024). These methods primarily target NISQ-era circuits and optimize structural metrics such as gate count, circuit depth, or routing overhead. In such settings, partial constructions admit meaningful intermediate cost estimates (e.g., remaining routing distance or local gate overhead), enabling effective guidance of the search before a circuit is fully specified.
@@ Acknowledgements
We thank the anonymous reviewers for their constructive feedback. This work was supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research under Contract No. DE-AC05-00OR22725 through the Accelerated Research in Quantum Computing Program MACH-Q project, the U.S. National Science Foundation CAREER Award No. CCF-2338773, and ExpandQISE Award No. OSI 2427020. GL was also supported by the Intel Rising Star Award.
@@ A.3.4. Data sets
  • ”family”: name of the QECC, string.
  • ”n”: number of physical qubits, integer.
  • ”k”: number of logical qubits, integer.
  • ”d”: code distance, integer.
  • ”logical_xs”: all logical $X$ operators, list of string.
  • ”logical_zs”: all logical $Z$ operators, list of string.
  • ”x_stabilizers,z_stabilizers”: all stabilizers, list of string.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 22–26, 2026; Pittsburgh, PA, USA.booktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’26), March 22–26, 2026, Pittsburgh, PA, USAisbn: 979-8-4007-2359-9/2026/03doi: 10.1145/3779212.3790123submissionid: V2fp0576ccs: Computer systems organization Quantum computingccs: Hardware Quantum error correction and fault toleranceccs: Software and its engineering Compilers
@@ 2.3. Monte Carlo Tree Search
    Selection: Starting from the initial state $S_{0}$, the algorithm traverses downward the tree by selecting child nodes that maximize the "upper confidence bound applied to trees" (UCT) (Kocsis and Szepesvári, 2006). For the $i$th child: $UCT_{i}=\hat{E_{i}}+c\sqrt{\frac{\ln N}{n_{i}}}$ where $\hat{E_{i}}$ is the expectation score of the $i$th child, $n_{i}$ is the number of times the child has been visited, $N$ is the total number of times the parent node has been visited, and $c$ is an exploration parameter, usually $\sqrt{2}$.
@@
    Expansion: The selection is continuously applied downward to the tree until either a termination or a not fully explored (it has non-child next states) node $S_{l}$. If $S_{l}$ is a termination state, then define $S_{e}:=S_{l}$ and proceed to the next step. Otherwise, we create a new child $S_{new}$ of $S_{l}$ if $\exists m\in mov(S_{l}),S_{l}\xrightarrow{m}S_{new}$ and $S_{new}$ is not already $S_{l}$’s child. We set $S_{e}:=S_{new}$. $S_{e}$ is referred to as the "expanded node".
@@
    Simulation: From $S_{e}$, a "roll-out" is performed to the termination. Specifically, we continuously select a random move starting from $S_{e}$ until we reach a termination state $S_{ter}$. An evaluation is applied to $eva(S_{ter})=E$.
@@ 5.1.1. QECCs and Decoders
We select multiple QECCs and decoders as the benchmarks, ranging across codes that encode single or multiple logical qubits. They include: rotated surface code (Bombín and Martin-Delgado, 2007; Horsman et al., 2012), color code (Bombin and Martin-Delgado, 2006; Rodriguez et al., 2024), surface code with defects (Kitaev, 2003), bivariate bicycle (BB) code (Bravyi et al., 2024b), hyperbolic surface code (Breuckmann and Terhal, 2016; Albuquerque et al., 2009), and hyperbolic color code (Delfosse, 2013; da Silva and Jr, 2018; Vuillot and Breuckmann, 2022). For the decoder, we include the minimum-weight perfect matching (MWPM) (Higgott and Gidney, 2025), BP-OSD (Roffe et al., 2020), and hypergraph unionfind (Berlekamp et al., 1978; Lovász, 1972) decoders.
@@ 5.3.2. Space-Time Resource Estimation and Reduction
We estimate absolute execution time using the IBM Brisbane device model, where a two-qubit gate costs approximately $600$ ns and a measurement takes approximately $4000$ ns (Zheng and Kanno, 2024). Single-qubit gate latency is negligible at this scale. The execution time of one syndrome-measurement round is therefore computed as: $T_{\text{round}}=\#\text{depth}\times T_{2Q}+T_{\text{meas}}$, and the corresponding space-time volume is given by $\text{Space-Time Volume}=T_{\text{round}}\times\#\text{physical qubits}$. Table 3 reports both the absolute execution time and the resulting space–time volume. While AlphaSyndrome sometimes incurs a higher depth in terms of two-qubit gates, this does not significantly increase—and in some cases even reduces—the absolute execution time required to achieve a given logical error rate. This effect arises from two factors. First, to reach comparable reliability, lowest-depth baseline schedules require larger code distances, whose minimal-depth schedules themselves have larger circuit depths. Second, syndrome-measurement rounds are dominated by the ancilla measurement latency, which is substantially longer than two-qubit gate latency (this holds for most hardware platforms); as a result, moderate increases in two-qubit gate depth are largely amortized by the fixed measurement cost.
@@ 5.5. Cross-Decoder Evaluation
Table 4 shows that AlphaSyndrome tailors its schedule to the target decoder. When tested with BP‑OSD, schedules compiled using BP‑OSD reduce the overall logical error rate by 25.4% on average compared to schedules compiled with unionfind, winning in 7/8 instances and peaking at 92.4% reduction on $[[7,1,3]]$. Symmetrically, when tested with unionfind, schedules compiled using unionfind reduce errors by 34.3% on average over BP‑OSD‑compiled schedules, again in 8/8 instances, with a peak of 55.53%. Overall, cross‑testing confirms that using a schedule compiled for a different decoder generally degrades performance, indicating that AlphaSyndrome indeed optimizes towards decoder behavior.
@@ 6. Related Work and Discussion
Syndrome Measurement Scheduling The syndrome measurement circuit scheduling problem for QEC codes plays an essential role in realizing fault-tolerant quantum computation. Some prior works on QEC discuss circuit scheduling for different QEC codes. For example,  (Horsman et al., 2012; Tomita and Svore, 2014; Dennis et al., 2002) examined circuit scheduling for surface code. In  (Tomita and Svore, 2014), the authors propose executing the gate following an S- or Z-shape ordering to minimize the effect of hook error. However, beyond surface code, this problem has not been well optimized. For instance, (Lacroix et al., 2025) adopts a trivial ordering using the qubit index for measuring syndromes in their color code experiments. Similarly,  (Pecorari et al., 2025) also adopts the trivial ordering for their hypergraph product code experiment, though this is a result of an analysis of hook error on effective code distance on hypergraph product code (Manes and Claes, 2025). Furthermore,  (Bravyi et al., 2024a) designed the syndrome extraction circuit to have minimal circuit depth for the BB code. However, these prior works only focus on a specific family of code, and their manual design approach cannot be automatically extended to other codes. Moreover, some works  (Lacroix et al., 2025; Pecorari et al., 2025) ignore the influence of hook error and focus only on the overall depth of the circuit. Even for works that consider hook error  (Tomita and Svore, 2014; Manes and Claes, 2025), their analysis remains at the circuit level without evaluating the logical error rate. Moreover, the approximation in the decoder and the non-uniform errors are not yet considered, leading to performance degradation. In contrast, the AlphaSyndrome proposed in this paper can automatically take all these factors into consideration and yield a syndrome measurement circuit implementation with much lower logical error rates.
QEC Mapping and Gadget Synthesis Several prior works develop synthesis techniques for other components of the QEC stack. QECC-Synth (Yin et al., 2025) employs MaxSAT to synthesize QEC layouts on sparse hardware architectures by optimizing qubit placement, routing structures, and ancilla-bridge usage, while Chiew et al. (Chiew et al., 2024) study SWAP-based embedding techniques that map fault-tolerant circuits onto constrained-connectivity hardware while preserving fault-tolerance guarantees. These approaches operate at the mapping and layout layer of the stack and do not modify the internal execution order of stabilizer measurements.
@@
MCTS-Based Quantum Synthesis and Routing MCTS has been applied in several quantum synthesis and optimization contexts, including AlphaRouter (Tang et al., 2024) and other MCTS- or reinforcement-learning–based approaches for routing and unitary synthesis (Valcarce et al., 2025; Rietsch et al., 2024). These methods primarily target NISQ-era circuits and optimize structural metrics such as gate count, circuit depth, or routing overhead. In such settings, partial constructions admit meaningful intermediate cost estimates (e.g., remaining routing distance or local gate overhead), enabling effective guidance of the search before a circuit is fully specified.
@@ Acknowledgements
We thank the anonymous reviewers for their constructive feedback. This work was supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research under Contract No. DE-AC05-00OR22725 through the Accelerated Research in Quantum Computing Program MACH-Q project, the U.S. National Science Foundation CAREER Award No. CCF-2338773, and ExpandQISE Award No. OSI 2427020.
@@ A.3.4. Data sets
  • "family": name of the QECC, string.
  • "n": number of physical qubits, integer.
  • "k": number of logical qubits, integer.
  • "d": code distance, integer.
  • "logical_xs": all logical $X$ operators, list of string.
  • "logical_zs": all logical $Z$ operators, list of string.
  • "x_stabilizers,z_stabilizers": all stabilizers, list of string.